Perfecto Pizza Company produces microwavable pizzas. The following accounts appeared in Perfectos ledger as of December 31.
Answer #1
Raw-Material Inventory Accounts Payable
Bal.1/1 $   21,200.00 $        2,500.00 Bal.1/1
$ 135,000.00 $ 120,200.00 $ 136,500.00 $    135,000.00
Bal. 12/31 $   36,000.00 $        1,000.00 Bal. 12/31
Work in Process Inventory Finished Goods Inventory
Bal.1/1 $   17,000.00 Bal.1/1 $   12,000.00
Direct Material $ 120,200.00 $ 718,200.00 $    710,200.00
Direct Labour $ 150,000.00 $ 718,200.00 Bal. 12/31 $   20,000.00
Manufacturing Overhead $ 450,000.00
Bal. 12/31 $   19,000.00
Manufacturing Overhead Cost of good Sold
$ 452,600.00 $ 450,000.00 $ 710,200.00
$ 710,200.00
Wages Payable Sales Revenue
$     2,000.00 Bal.1/1 $    809,900.00
$ 147,000.00 $ 150,000.00
$     5,000.00 Bal. 12/31
Accounts Receivables
Bal.1/1 $   11,100.00
$ 809,900.00 $ 806,000.00
Bal. 12/31 $   15,000.00

Accounts Payable 3 Bal.1/1 Bal 1/1 Raw Material Inventory 21200 =H4 36000 =B3+B4-B5 136500 2500 =G4+H5-H3 1000 Bal. 12/31 Bal